OpenStudy (anonymous):

What is the equation of a line that passes through the point (6,3) and is perpendicular to a line with a slope -3/2

hero (hero):

The general form of the equation of a line is y = mx + b correct?

OpenStudy (anonymous):

i dont know thats why i am asking for help

OpenStudy (anonymous):

i would say yes

OpenStudy (anonymous):

the reciprocal is the slope but what is b

hero (hero):

And you would be correct. The slope of the equation of the line is m = 2/3, right?

OpenStudy (anonymous):

yes

hero (hero):

So we can input m = 2/3 into the equation of the line as follows: y = (2/3)x + b Also, we are given a point on the line. (x,y) = (6,3) correct?

OpenStudy (anonymous):

so what is B ......i think im just slow or i have bad communication...........how do i find b what is it.

Nnesha (nnesha):

put x y value in to the equation

hero (hero):

3 = (2/3)(6) + b 3 = 2(2) + b 3 = 4 + b

hero (hero):

Finish solving for b

OpenStudy (anonymous):

what does that mean ..........it says it can only be -1 or 3 so y=2/3x-1 or y=2/3x+3

OpenStudy (anonymous):

HOW

hero (hero):

3 - 4 = b -1 = b

OpenStudy (anonymous):

WHERE ARE YOU GETTIN 3-4

hero (hero):

You have to plug in the values then solve for b algebraically.

hero (hero):

y = mx + b <-You start with this 3 = (2/3)(6) + b <-After plugging in the known values you get this 3 = 2(2) + b <-After simplification you get this 3 = 4 + b 3 - 4 = b <-You get this after isolating b (subtracting 4 from both sides) -1 = b <-This is the result